OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6.

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6. y = -x + 4 y = 3x + 20 y = x y = -3x - 16

OpenStudy (anonymous):

HELLPPPP